# Noise Synthesis for Low-Light Image Denoising with Diffusion Models

### v1

@@ 4.3 Image Denoising Performance Comparison
As stated in Sec. 4.1, we compare the performance of denoising networks trained with data from different synthesis methods. We test two physics-based methods (the Poisson-Gaussian model and ELD [52]), and two recent learning-based methods (LRD [56] and NoiseFlow [2]) [^1] For ELD and LRD, we use pre-trained models available on the authors’ official websites. For NoiseFlow, we retrain it on the SID dataset.. Notably, our evaluation differs from ELD and LRD in two key ways:
  • ELD crops the image centers to a size of $512\times 512$ for PSNR/SSIM computation, whereas we compute them on the entire image at a resolution of $4240\times 2832$.
  • In LRD’s image denoising code, a precomputed dark shading frame, which is computed by averaging multiple dark frames as mentioned in Sec. 2, is subtracted from the noisy images before denoising [^2] See this part of the code.. We omit this step, as such data is unavailable in usual cases.
In the following results, we apply the same measurement setup for all methods.
Figure 4: Visual comparison of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. The test sample is from the SID test set. All results here are processed by the same gamma correction for visualization purposes. Our method produces clearer details and effectively addresses the low-frequency noise at the bottom of the image, whereas other methods produce purplish artifacts. (Best viewed when zoomed in)
Numerical Results. Table 3.3 presents quantitative comparisons between different methods. The upper part of the table shows results obtained without any post-processing, where the denoised images are directly evaluated using classic metrics. The lower part presents the results after post-processing. Since PSNR and SSIM metrics are sensitive to overall image illumination, even minor shifts can lead to significant variations in these values. We apply the illumination correction method from the ELD paper [52]  [^3] See this part of the code. to the denoised images of all methods before evaluating their PSNR and SSIM.
Our approach achieves the best results on the SID dataset compared to other data synthesis techniques. It also outperforms the model trained with real data pairs in most cases. On the ELD test set, our method shows the best performance at the ×$\times 100$ exposure ratio among all data synthesis approaches. However, our method achieves relatively lower PSNR/SSIM results than ELD at the ×$\times 200$ exposure ratio (though still outperforming the learning-based LRD method and comparable to the real paired data). This is because the ELD method generates noisy images by sampling noise parameters from a realistic distribution, thus covering a broader range of noise levels. As our diffusion model was not trained with data from the ×$\times 200$ exposure ratio (as this ratio is not present in the SID dataset), its generalization capability is somewhat limited. However, as we will show in the qualitative analysis, our method still produces denoised images with less noise and better detail preservation compared to other methods. The primary discrepancy is a color shift, which is a potential area for future improvement.
@@
Qualitative Results. We compare denoising networks on an example from the SID test set in Fig. 4. While other noise synthesis methods struggle with low-frequency noise that appears at the bottom of the image, our method addresses this effectively. Visual comparisons on the ELD test set in Fig. 1 show that our method recovers the text details more clearly than other methods. Additional examples are given in the appendix.
@@ 4.4 Discussions
In this section, we assess the importance of various components in our diffusion model by analyzing the characteristics of the generated noise and the denoising performance of networks trained on data generated by ablated models
Ratio Ours (Limited data) Ours
@@
Table 2: Ablation study on the influence of training-data diversity for denoising. We show the PSNR/SSIM results on the SID test set.
Influence of Training-Data Diversity. The number of scenes per camera setting in the SID training set is uneven, with some settings having more than ten scenes and some only one (e.g., 28 scenes for ISO 250 and 1 for ISO 2500). To evaluate the influence of training-data diversity, we train a denoising network on synthetic data using the same number of scenes per camera setting as in the SID training set. The results, reported as Ours (Limited data) in Table 2, are unsurprisingly close to those obtained when training on paired real data, as shown in Table 3.3.
To leverage the full potential of our diffusion model, we generate more synthetic data by randomly sampling scenes from the SID dataset, regardless of their camera settings. This sampling strategy improves image denoising performance, as shown in Table 2. Visual comparisons in the appendix further support that it improves the results for under-represented camera settings. Thus, our method offers a promising approach for improving denoising performance when real-world data is scarce.
Sampling Method. The generation speed of our method is constrained by the many steps required by DDPM sampling. Therefore, in the following ablation studies, we use DDIM sampling [45] as an alternative, which uses a non-Markovian framework to achieve faster generation with fewer steps. We first validate that DDIM could achieve satisfactory performance with a minor drop compared to DDPM. We train a denoising network with synthetic data generated using 100-step DDIM sampling. As shown in Fig. 5 (d), DDIM generates noise resembling real noise, with slight variance inaccuracy. Table 3 (C) and (D) show that while DDIM introduces a performance trade-off, the decrease is acceptable.
Figure 6: Comparison between real and synthetic data. The samples in the two rows are from the SID Sony test set (not used for diffusion training) under the following settings: ISO 6,400 with a ratio of 300; and ISO 200 with a ratio of 250. Different noise values are represented by varying colors, with the color bar displayed on the right. Our method generates realistic noise that closely resembles real noise.
@@
Table 3: Ablation study on network architectures and diffusion sampling methods. We present the PSNR/SSIM results of denoising networks trained with different synthetic data sources generated by various methods on the SID test set. We also report the KL divergence (in italic) between the noise distributions produced by these methods and the real noise distribution.
@@
Effects of Two-Branch Network Architecture. We investigate the effects of our proposed two-branch network architecture by conducting an ablation study, which adopts a sole UNet architecture for the diffusion model. This UNet-only architecture takes the clean image $\mathbf{x}$ and the output of the last step $\tilde{\mathbf{n}}_{t}$ as inputs; it then outputs the result of the current step $\tilde{\mathbf{n}}_{t-1}$. To assess performance differences, we generate synthetic data from clean images in the SID training set using both architectures with DDIM sampling for efficiency. We then compare the performance of two denoising networks trained on these respective synthetic datasets, as shown in Table 3. Comparing (A) and (C) shows that the full network achieves higher PSNR/SSIM. Additionally, the full model’s noise has a smaller KL divergence from real noise than the UNet-only model. We also condition both architectures on the same clean images from the SID test set and visualize the generated results. Examples are shown in Fig. 5 (b) and (e). Note that our diffusion network generates patches of $512\times 512$ size. To visualize the full image, we piece patches together. The noise patches produced by the UNet-only architecture exhibit a checkerboard pattern; this indicates inconsistent mean values among patches. This phenomenon suggests that the UNet-only architecture generates noise images with unstable mean values.
@@
Effects of Positional Encoding. We further investigate the effects of positional encoding. We conduct the following experiment to assess its effects: During inference, we modify the generation process to $\tilde{\mathbf{n}}_{t}=f_{\theta}(\tilde{\mathbf{n}}_{t-1},t,\mathbf{x},\mathbf{0},\mathbf{z})$, where $\mathbf{0}$ denotes a black image filled with zeros, signifying that the diffusion model is conditioned on a zero map instead of on the exact pixel coordinates. We compare KL divergence of noise generated by models with and without positional encoding. Table 3 (B) and (C) show that the denoising network trained with data generated with positional encoding yields higher PSNR and SSIM, and the KL divergence increases when positional encoding is not used. The generated results are depicted in Fig. 5 (c). By comparing it with the full results in Fig. 5 (e), it can be observed that the generated noises do not contain the low-frequency noise in the bottom; this indicates the effectiveness of positional encoding in modeling the position-related noises.
@@
Noise Visualization. In Fig. 6, we visualize noise images generated by our model. By comparing real and generated noise, we demonstrate that our diffusion model accurately captures realistic noise patterns. In the first sample, the real noise image exhibits low-frequency noise at the bottom. Our model successfully reproduces this, as the generated noise also shows low-frequency noise at the bottom. Failure to capture this pattern would result in a purplish offset in the final denoised images. As shown in Fig. 4, our denoising network effectively handles low-frequency noise, unlike other methods, which produce purplish artifacts at the bottom.
@@
5 Conclusion and Limitations
Diffusion models have been widely applied for generating realistic images, but their use for generating real-world noise remains unexplored. Nonetheless, We find that a direct application of diffusion models is insufficient for realistic noise synthesis. Our tailored diffusion framework enables generating synthetic noise that closely matches real-world noise, thus outperforming existing noise synthesis methods. Moreover, our analysis of the generated data provides valuable insights into using diffusion models for real-world noise synthesis. Despite the promising results, our method also shows some limitations in generalization, especially in generalizing to unseen camera settings. Future work could improve generalization by integrating physics-based noise modeling with our diffusion-based approach. Generation speed is also a constraint; while DDIM sampling offers some acceleration, it still incorrectly maps the noise variance. Consequently, faster and more accurate approaches could be explored.
Appendix
@@ 5 Conclusion and Limitations
  • Preliminaries: Diffusion Models $\rightarrow$ Sec. A
  • Proof for Variance Preserving Noise Schedule $\rightarrow$ Sec. B
  • Implementation Details $\rightarrow$ Sec. C
    • Network Architecture $\rightarrow$ Sec. C.1
    • Datasets $\rightarrow$ Sec. C.2
    • Training Configurations $\rightarrow$ Sec. C.3
    • Clean-Noisy Pair Generation $\rightarrow$ Sec. C.4
  • Evaluation Protocol: Noise Mean and Variance $\rightarrow$ Sec. D
  • More Discussions $\rightarrow$ Sec. E
    • Influence of Training-Data Diversity $\rightarrow$ Sec. E.1
    • Sampling Method $\rightarrow$ Sec. E.2
    • Noise Decomposition $\rightarrow$ Sec. E.3
    • Noise Histogram $\rightarrow$ Sec. E.4
  • More Visual Comparisons $\rightarrow$ Sec. F
@@
Appendix A Preliminaries: Diffusion Models
@@
Appendix B Proof for Variance Preserving Noise Schedule
In this section, we present the proof of Proposition 1 discussed in the main paper.
@@ Appendix B Proof for Variance Preserving Noise Schedule
Given a signal $u\sim\mathcal{N}(\lambda,\sigma_{2}^{2})$ and its measurement $v=u+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma_{1}^{2})$, the conditional distribution of $p(U|V)$ is
@@
As previously discussed, in our context, the target distribution $U$ that we aim to recover in the reverse process is the low-light noise distribution. For simplicity, we approximate it as a Gaussian distribution without sacrificing generality. Since the primary components of low-light noise (such as shot noise, read noise, and banding patterns) are either inherently modeled as Poisson distributions or can be approximated by Gaussian distributions [52]. Moreover, with a sufficient photon count, a Poisson distribution can be approximated by a Gaussian, as a consequence of the central limit theorem.
Appendix C Implementation Details

### v2

@@ 4.4 Image Denoising Performance Comparison
As stated in Sec. 4.1, we compare the performance of denoising networks trained with data from different synthesis methods. We evaluate three physics-based methods (the Poisson-Gaussian model, ELD [9] and PMN [60]), and three recent learning-based methods (Starlight[11], LRD [12] and NoiseFlow [13]) [^2] For ELD, PMN and LRD, we use pre-trained models available on the authors’ official websites. For NoiseFlow, we retrain it on the SID dataset. The results of Starlight are taken from the LLD paper[44].. We also compare our method with a denoising network trained with real data pairs from the SID Sony training set.
Among these methods, PMN and LRD apply dark shading correction (DSC) for noise calibration, where a precomputed dark shading frame—obtained by averaging multiple dark frames, as described in Sec. 2—is subtracted from the input noisy images during both denoising training and inference. In contrast, the Poisson-Gaussian model, ELD, and NoiseFlow do not incorporate this correction.
For a fair comparison, we train two versions of our denoising network. The first, NoiseDiff, is trained directly on our synthetic data without DSC. The second, NoiseDiff*, is trained on our synthetic data processed with DSC by subtracting the dark shading.
Additionally, NoiseDiff* employs shot noise augmentation (SNA) proposed in PMN. This method relies on the linear behavior of shot noise, by artificially scaling the ground-truth image and adding Poisson noise accordingly to the noisy images. Specifically, the clean image $I$ is first scaled by a random factor $\delta$, resulting in a clean signal increment $\Delta=(1-\delta)I$. A noisy signal increment is then sampled from a Poisson distribution $\mathcal{P}(\Delta)$ and added to the noisy image, in line with the image formation process described in Sec. 2. In practice, SNA enhances the diversity of shot noise in our synthetic data. In LRD, shot noise is sampled from a Poisson distribution on the fly during training, making this process equivalent to SNA. This equivalence justifies the comparison with our method augmented with SNA. For more details about DSC and SNA, please refer to the appendix.
It is also worth noting that, in addition to augmented data, PMN also uses real data pairs from the SID Sony training set during training of the denoising network. In contrast, all other data synthetic methods, including our NoiseDiff and NoiseDiff*, use only synthetic data.
Numerical Results. Table 3.3 presents quantitative comparisons between different methods. The middle part compares NoiseDiff without calibration to other calibration-free methods, while the right part evaluates NoiseDiff* against methods that incorporate dark shading data for calibration.
As shown in Table 3.3, our NoiseDiff achieves the best results on the SID dataset among calibration-free data synthesis methods and outperforms models trained with real data pairs in most cases. Our method also demonstrates competitive PSNR/SSIM performance on the ELD dataset.
Dark shading calibration eases the training of the denoising network, as it centers noise on zero, by removing the inhomogeneous bias in dark frames. Therefore, DSC calibrated denoising networks tend to perform better than previous ones. Among them, our NoiseDiff* achieves the best performance in most cases. For example, on SID ×$\times 100$, it surpasses PMN, the second-best approach, by 0.45 dB in PSNR.
@@
Qualitative Results. We compare denoising performance on two examples from the SID and ELD test sets in Fig. 4. In the zoomed-in area of the SID example, our method produces cleaner results, while other methods leave some residual noise. The second example from the ELD test set demonstrates that our method recovers text details more clearly and with fewer artifacts. Additional examples in the appendix show that, while other calibration-free noise synthesis methods struggle with low-frequency noise appearing at the bottom of the image, our method effectively addresses this issue even without calibration data.
@@ 4.5 Discussions
In this section, we assess the importance of various components in our diffusion model by analyzing the characteristics of the generated noise and the denoising performance of networks trained on data generated by ablated models. In the following ablation experiments, we exclude dark shading correction, shot noise augmentation, and illumination correction when assessing denoising performance. This ensures that the results reflect the direct network output, making it easier to isolate and evaluate the effects of our main contributions.
TABLE III: Ablation study on the influence of training-data diversity for denoising. We show the PSNR/SSIM results on the SID test set.
Ratio NoiseDiff (Limited data) NoiseDiff
@@
Influence of Training-Data Diversity. The number of scenes per camera setting in the SID training set is uneven, with some settings having more than ten scenes and some only one (e.g., 28 scenes for ISO 250 and 1 for ISO 2500). To evaluate the influence of training-data diversity, we train a denoising network on synthetic data using the same number of scenes per camera setting as in the SID training set. The result is reported as NoiseDiff (Limited data).
To leverage the full potential of our diffusion model, we generate more synthetic data by randomly sampling scenes from the SID dataset, regardless of their camera settings. This sampling strategy improves image denoising performance, as shown in Table III. Visual comparisons in the appendix further support that it improves the results for under-represented camera settings. Thus, our method offers a promising approach for improving denoising performance when real-world data is scarce.
Sampling Method. The generation speed of our method is constrained by the many steps required by DDPM sampling. We tested our model on a server with dual AMD EPYC 7543 32-core CPUs (128 threads total) and an NVIDIA A100-SXM4-80GB GPU. The average time for generating a $512\times 512$ patch using DDPM (1000 steps) is $43.6~{}s$, excluding IO time.
Therefore, in the following ablation studies, we use DDIM sampling [61] as an alternative, which uses a non-Markovian framework to achieve faster generation with fewer steps. We first validate that DDIM could achieve satisfactory performance with a minor drop compared to DDPM. We train a denoising network with synthetic data generated using 100-step DDIM sampling. As shown in Fig. 5 (d), DDIM generates noise resembling real noise, with slight variance inaccuracy. Table IV (C) and (D) show that while DDIM introduces a performance trade-off, the decrease is acceptable.
Figure 6: Comparison between real and synthetic data. The samples in the three rows are from the SID Sony test set (not used for diffusion training) under the following settings: ISO 6,400 with a ratio of 300; ISO 200 with a ratio of 250; and ISO 12800 with a ratio of 100. Different noise values are represented by varying colors, with the color bar displayed on the right. Our method generates realistic noise that closely resembles real noise.
TABLE IV: Ablation study on network architectures and diffusion sampling methods. We present the PSNR/SSIM results of denoising networks trained with different synthetic data sources generated by various methods on the SID test set. We also report the KL divergence (in italic) between the noise distributions produced by these methods and the real noise distribution.
@@
Effects of Two-Branch Network Architecture. We investigate the effects of our proposed two-branch network architecture by conducting an ablation study, which adopts a sole UNet architecture for the diffusion model. This UNet-only architecture takes the clean image $\mathbf{x}$ and the output of the last step $\tilde{\mathbf{n}}_{t}$ as inputs; it then outputs the result of the current step $\tilde{\mathbf{n}}_{t-1}$.
@@
To assess performance differences, we train this UNet-only model from scratch using the same training setup as the full network. We then generate synthetic data from clean images in the SID training set using both the full network and the UNet-only network, employing DDIM sampling for efficiency. We compare the performance of two denoising networks trained on these respective synthetic datasets, as shown in Table IV. Comparing (A) and (C) shows that the full network achieves higher PSNR/SSIM. Additionally, the full model’s noise has a smaller KL divergence from real noise than the UNet-only model. Note that the KLD values for our full model in Table IV differ slightly from those reported in Table II. This is because, for efficient ablation studies, we sample a subset of the data instead of using the full dataset. In contrast, Table II reports the KLD calculated over the entire dataset.
We also condition both architectures on the same clean images from the SID test set and visualize the generated results. Examples are shown in Fig. 5 (b) and (e). Note that our diffusion network generates patches of $512\times 512$ size. To visualize the full image, we piece patches together. The noise patches produced by the UNet-only architecture exhibit a checkerboard pattern; this indicates inconsistent mean values among patches. This phenomenon suggests that the UNet-only architecture generates noise images with unstable mean values.
@@
Effects of Positional Encoding. We further investigate the effects of positional encoding. We conduct the following experiment to assess its effects: During inference, we modify the generation process to $\tilde{\mathbf{n}}_{t}=f_{\theta}(\tilde{\mathbf{n}}_{t-1},t,\mathbf{x},\mathbf{0},\mathbf{z})$, where $\mathbf{0}$ denotes a black image filled with zeros, signifying that the diffusion model is conditioned on a zero map instead of on the exact pixel coordinates. We compare KL divergence of noise generated by models with and without positional encoding. Table IV (B) and (C) show that the denoising network trained with data generated with positional encoding yields higher PSNR and SSIM, and the KL divergence increases when positional encoding is not used. The generated results are depicted in Fig. 5 (c). By comparing it with the full results in Fig. 5 (e), it can be observed that the generated noises do not contain the low-frequency noise in the bottom; this indicates the effectiveness of positional encoding in modeling the position-related noises.
@@
Noise Visualization. In Fig. 6, we visualize noise images generated by our model. By comparing real and generated noise, we demonstrate that our diffusion model accurately captures realistic noise patterns. In the first sample, the real noise image exhibits low-frequency noise at the bottom. Our model successfully reproduces this, as the generated noise also shows low-frequency noise at the bottom. Failure to capture this pattern would result in a purplish offset in the final denoised images. As shown in some examples in the appendix, our denoising network effectively handles low-frequency noise, unlike other methods, which produce purplish artifacts at the bottom.
One can observe that in our generated noise images, the banding noise is visually less prominent compared to the real data. In real noise, the banding pattern appears as long horizontal lines, while in our synthetic noise images, the bands are shorter and less continuous. This may be attributed to the limited receptive field of our network and the inherent randomness of the diffusion sampling process. Nonetheless, as demonstrated by our qualitative denoising results (e.g., Fig. 4), our denoiser effectively removes banding noise, while competing methods such as LRD and ELD leave residual artifacts. This indicates that the shorter synthetic banding pattern in our data is sufficient for training a denoiser capable of handling the more pronounced banding pattern observed in real data.
@@
5 Limitations
Despite the promising results, our method has some limitations. As noted in the discussion, the banding noise in our generated noise images is visually less prominent compared to the real data. The generation speed is also relatively slow due to the many steps required by the DDPM sampling process. Although DDIM sampling offers some acceleration, it does not map the noise variance as accurately. Consequently, faster and more accurate approaches remain an area for further exploration. Additionally, our method still relies on a certain number of paired clean-noisy images for data synthesis training. Investigating unsupervised or self-supervised approaches could be a promising direction for future work.
6 Conclusion
Diffusion models have been widely applied for generating realistic images, but their use for generating real-world noise remains unexplored. Nonetheless, we find that a direct application of diffusion models is insufficient for realistic noise synthesis. Our tailored diffusion framework—featuring a two-branch architecture, pixel positional encoding, and carefully chosen noise schedules—enables the generation of synthetic noise that closely matches real-world noise, outperforming existing noise synthesis methods. Moreover, our analysis of the generated data provides valuable insights into leveraging diffusion models for real-world noise synthesis.
Appendix: Dark Noise Diffusion Noise Synthesis for Low-Light Image Denoising
@@ 6 Conclusion
  • Preliminaries: Diffusion Models $\rightarrow$ Sec. 7
  • Proof for Variance Preserving Noise Schedule $\rightarrow$ Sec. 8.1
  • Implementation Details $\rightarrow$ Sec. 9
    • Network Architecture $\rightarrow$ Sec. 9.1
    • Datasets $\rightarrow$ Sec. 9.2
    • Training Configurations $\rightarrow$ Sec. 9.3
    • Clean-Noisy Pair Generation $\rightarrow$ Sec. 9.4
  • Evaluation Protocol: Noise Mean and Variance $\rightarrow$ Sec. 10
  • More Discussions $\rightarrow$ Sec. 11
    • Influence of Training-Data Diversity $\rightarrow$ Sec. 11.1
    • Sampling Method $\rightarrow$ Sec. 11.2
    • Noise Decomposition $\rightarrow$ Sec. 11.3
    • Noise Histogram $\rightarrow$ Sec. 11.4
    • Influenc of Dark Shading Correction and Shot Noise Augmentation $\rightarrow$ Sec. 11.5
  • More Visual Comparisons $\rightarrow$ Sec. 12
@@
7 Preliminaries: Diffusion Models
@@
8 More Details about Variance Preserving Noise Schedule
In this section, we first provide the proof of Proposition 1 discussed in the main paper. We then present an additional experiment on 1D Tukey-Lambda distributions to further validate the importance of noise schedules.
8.1 Proof for Proposition 1
@@
Given a signal $u\sim\mathcal{N}(\lambda,\sigma_{2}^{2})$ and its measurement $v=u+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma_{1}^{2})$, the conditional distribution $p(U|V)$ can be derived using properties of Gaussian distributions and Bayesian inference [62]. It is given by
@@
As discussed previously, in our context, the target distribution $U$ that we aim to recover in the reverse process is the low-light noise distribution. For simplicity, we approximate it as a Gaussian distribution without sacrificing generality. Since the primary components of low-light noise (such as shot noise, read noise, and banding patterns) are either inherently modeled as Poisson distributions or can be approximated by Gaussian distributions [9]. Moreover, with a sufficient photon count, a Poisson distribution can be approximated by a Gaussian, as a consequence of the central limit theorem.
More generally, as mentioned in the main paper, the same theory still holds for the Gaussian Mixture Models (GMMs). Assume the target distribution follows a GMM: $p(U)\sim\sum_{k=1}^{K}w_{k}\mathcal{N}(U;\mu_{k},{\tau_{k}}^{2})$, where $w_{k}$, $\mu_{k}$, ${\tau_{k}}^{2}$ are the mixture weights, means, and variances of the $k$-th Gaussian component, we have $p(U|V=v)=\sum_{k=1}^{K}\pi_{k}(v)\mathcal{N}\left(U;\mu_{k}+\frac{\tau_{k}^{2}}{\tau_{k}^{2}+\sigma_{1}^{2}}(v-\mu_{k}),\frac{{\tau_{k}}^{2}\sigma_{1}^{2}}{{\tau_{k}}^{2}+\sigma_{1}^{2}}\right)$, where $\pi_{k}(v)$ is the posterior weight. Thus the variance of the MMSE solution is $\sum_{k=1}^{K}\pi_{k}(v)\left(\frac{\tau_{k}^{2}}{\tau_{k}^{2}+\sigma_{1}^{2}}\right)^{2}\left(\sigma_{1}^{2}+\text{Var}(u)\right)$.
Figure S1: Experiments on seven TL distributions with different $\sigma$, where the x axis denotes the scale parameter $\sigma$ of these seven distributions, and the y axis shows the ratio of standard deviation between generated and GT samples. As can be observed, sigmoid3 (and sigmoid2) produces standard deviations closer to the ground truth, while cosine yields consistently lower stds
8.2 Experiment on 1D Tukey-Lambda Distributions
We extend the experiment shown in Fig. 3 to a set of 1D Tukey-Lambda distributions, which have been shown in the ELD paper to closely approximate the distribution of read noise. Specifically, we set the target distribution to Tukey-Lambda distributions with fixed shape ($\lambda=0.1$) and location ($\mu=0$) parameters while varying the scale parameter $\sigma$. We train a new 1D diffusion network to model seven different predefined distributions.
As shown in Fig. S1, for different Tukey-Lambda distributions, sigmoid3 and sigmoid2 produce standard deviations closer to the ground truth, whereas cosine consistently yields lower standard deviations. This further validates our design.
Additionally, as the scale parameter $\sigma$ increases, the performance gap between different noise schedules narrows. This is because, as indicated in Eq.16, when the signal variance $\sigma_{2}^{2}$ increases, the influence of the added Gaussian noise variance $\sigma_{1}^{2}$ diminishes.
[…]
9 Implementation Details